# Quantum dispersionless KdV hierarchy revisited

### v2

@@ 2.2. Quantum Hamiltonian structure
where $f,g\in\mathcal{A}[\hbar]$ are arbitrary lifts of $F$ and $G$, respectively. This Lie bracket defines a deformation quantization of the Hamiltonian structure (2.6) in the sense that
@@
In particular, we have proved Theorem 1.1 and Proposition 1.4 as special cases of Theorem 2.7 for $N=1$ and $\eta=1$. As an application of the above formula, let us prove the following interesting fact which will be used later.
@@
The fact that $a_{(0)}=0$ implies that, for any $g\in\mathcal{A}$, $\varphi(f)_{(0)}\varphi(g)=0$. Then by taking the $\hbar$-coefficient of $\varphi(f)_{(0)}\varphi(g)$, we arrive at
@@
here we denote by $\eta_{\alpha\beta}$ the inverse of $\eta^{\alpha\beta}$. Due to (2.12), we see that $A=0$, hence we conclude that
@@
which implies that actually $\varphi(f_{0})_{(0)}=0$. By passing to $(f-f_{0})/\hbar$, the same argument shows that $f_{1}$ is of the same form as $f_{0}$. In this way, we prove that actually each $f_{n}$ has the same form of $f_{0}$, namely, there exist polynomials $B_{\alpha}(\hbar),C(\hbar)\in\mathbb{C}[\hbar]$ such that
@@ 3. Weyl quantization of dispersionless KdV hierarchy
$[H_{2}^{c},H_{3}^{c}]=\int\frac{\hbar}{4}v^{4}v_{x}+\frac{\hbar^{2}}{12}(3vv_{x}v_{xx}+v^{2}v_{xxx})+\frac{\hbar^{3}}{720}v^{(5)}\neq 0.$
Therefore, the quantization problem is to find a quantum correction $H_{n}$ of $H_{n}^{c}$ such that
$[H_{n},H_{m}]=0,\quad\lim_{\hbar\to 0}H_{n}=H_{n}^{c}.$
@@ 3.1. Construction of quantum Hamiltonians
From now on, let us denote by $V$ the vertex algebra of rank $=1$. It is generated by the field
$Y(b,z)=\sum_{n\in\mathbb{Z}}b_{(n)}z^{-n-1},\quad\left[b_{(n)},b_{(m)}\right]=\hbar\,n\,\delta_{m+n}.$
Our idea to construct the quantum Hamiltonians is as follows. The ring $\mathcal{A}$ of differential polynomials is a commutative algebra, and we can view the normal order product on $V$ as a deformation of this algebra structure, in the sense that
@@
here recall that $:ab:=a_{(-1)}b$ for $a,b\in V$. The normal product is neither commutative nor associative, and we can use the idea of Weyl quantization in quantum mechanics to quantize the Hamiltonians $H_{n}^{c}$. Roughly speaking, in Weyl quantization, a quantum Hamiltonian is obtained from a classical Hamiltonian by averaging all the possible ways to compose all the quantum operators appeared in Hamiltonian. For example, let $(x,p)$ be a pair of conjugated variables, then we have the following Weyl quantization:
$(xp)^=\frac{1}{2}(\hat{x}\hat{p}+\hat{p}\hat{x}),\quad(xp^{2})^=\frac{1}{3}(\hat{x}\hat{p}\hat{p}+\hat{p}\hat{x}\hat{p}+\hat{p}\hat{p}\hat{x}).$
@@
To give a precise definition of the quantum Hamiltonians, let us define the set $BT(n)$ of labeled full binary rooted tress with $n$ leaves for $n\geq 1$ recursively. Each element in $BT(n)$ will be a full binary tree whose nodes, except for the root node, are labeled either by $L$ (for left) or $R$ (for right). In addition, $n$ leaves of a tree in $BT(n)$ are labeled by $1,\dots,n$. $BT(1)$ is just the set of a single root node and $BT(2)$ is the set of the unique full binary tree with two leaves labeled by $L,1$ and $R,2$.
RootL,1R,2
@@
For each tree $T\in BT(n)$ and for a choice of its leaf with label $i\in\{1,\dots,n\}$, let us define a new tree $T_{i}$ as follows. Firstly we make this leaf with label $i$ an internal node by adding to it 2 leaves, and hence obtain a full binary tree with $n+1$ leaves. Then the label ‘$i$’ of this node is removed. For leaves originally labeled by $11,\dots,i-$ 1 their labels remain unchanged, and for leaves originally labeled by $i+1,\dots,n$ their labels are added by one. Finally, the newly added leaves are labeled as $L,i$ and $R,i+1$. We then define
@@
From the construction, it is easy to see that $|BT(n)|=(n-1)!$. As an example, elements in $BT(3)$ is presented in Fig. 2.
RootLL,1R,2R,3
RootL,1RL,2R,3
@@
Let us proceed to associate a state to each tree in $BT(n)$. For each tree $T\in BT(n)$, we first put at every leaf the state $b\in V$, and then we recursively put at every other nodes the state obtained by taking the normal order product of states associated with its left child and right child. More explicitly, given a node which is not a leaf, assume that we have put $f\in V$ at its left child and $g\in V$ at its right child, then this node is associated with the state $:fg:\in V$. We denote by $A_{T}\in V$ the state associated with the root of $T$. For example, for $T_{1},T_{2}\in BT(3)$ as presented in Fig. 2, we have
@@
By using Corollary 2.6, it is easy to see that
@@
hence let us define the quantum Hamiltonian densities
$h_{n}=\frac{1}{(n+1)!}\varphi^{-1}(Z_{n+1}),\quad n\geq 0.$
Alternatively, we can equivalently define $h_{n}$ by the recursion relation
@@
It is obvious that $h_{0}=v$ and $h_{1}=\frac{v^{2}}{2}$. Let us compute $h_{2}$. First we see
$:b:bb::=b_{(-1)}b_{(-1)}b_{(-1)}|0\rangle,$
and by using the quasi-associativity (2.2) of normal order product, we have
$::bb:b:=:b:bb::+:\left(\int_{0}^{T}d\lambda\,b\right)[b_{\lambda}b]:=:b:bb::+\hbar T^{2}b,$
@@
$[b_{\lambda}b]=\sum_{n\geq 0}\frac{\lambda^{n}}{n!}b_{(n)}b_{(-1)}|0\rangle=\hbar\lambda|0\rangle.$
@@
Let us compute $h_{3}$. We begin by enumerate elements in $BT(4)$. Denote by $R$ and $S$ the first and the second tree presented in Fig 2, then by definition we arrive at
@@
These trees are presented in the following diagrams.
RootLLL,1R,2R,3R,4
RootLL,1RL,2R,3R,4
RootLL,1R,2RL,3R,4
@@
Let us compute the contribution from each graph. We start from
@@
$:b^{2}b^{2}:-b^{4}=2\left(\int_{0}^{T}d\lambda\,b\right)[b_{\lambda}b^{2}]=2\hbar:T^{2}bb:,\quad b^{4}=\,:b:b:bb:::,$
@@
$[b_{\lambda}b^{2}]=:[b_{\lambda}b]b:+:b[b_{\lambda}b]:+\int_{0}^{\lambda}[[b_{\lambda}b]_{\mu}b]d\mu=2\hbar\lambda b.$
From the above computation we see that
@@
$\left(-2\hbar\frac{\partial}{\partial\hbar}+\sum_{k\geq 0}kv^{(k)}\frac{\partial}{\partial v^{(k)}}\right)h_{n}$ $=0.$
@@
Note that by definition, each $Z_{n}$ is of conformal degree $n$, where we define each $b_{(n)}$ to be of conformal degree $-n$. This conformal degree of $V$ induces a gradation on $\mathcal{A}[\hbar]$ via $\varphi$, with $v^{(k)}$ to be of degree $k+1$, hence we prove the first identity.
@@
This gradation induces a gradation on $V$, where $\deg_{\partial_{x}}b_{(n)}=-n-1$. For $f\in\mathcal{A}[\hbar]$ of degree $d$, it is easy to see, using Proposition 2.1 that
@@
Then we can prove by induction, using the recursion (3.1), that each $h_{n}$ is of differential degree 00. The lemma is proved. ∎
@@ 3.2. Another expression for quantum Hamiltonians
As we have already seen, it is not easy to compute each $h_{n}$ using the definition. In what follows, we give another expression of $h_{n}$, which makes it easy to compute $h_{n}$, and more importantly, enables us to prove various properties of quantum Hamiltonians. From now on, for simplicity, we will set $\hbar=1$, and the dependence of $\hbar$ may be recovered from homogeneous conditions.
@@
This recursion relation makes it easy to compute each $g_{n}$, and in particular, it follows that $g_{n}$ are given by Schur polynomials. Recall that for a partition $\lambda=(\lambda_{1},\dots,\lambda_{\ell(\lambda)})$, we associate it with a differential polynomial $S_{\lambda}\in\mathcal{A}$, called the Schur polynomial associated with the partition $\lambda$, by
@@
the generating function of $g_{n}$, then we see that this function is uniquely determined by the equation
@@
Then it suffices to show that
$G(z)=\frac{W}{z}-\frac{1}{z},\quad W=\exp\left(\sum_{k\geq 0}\frac{v^{(k)}}{(k+1)!}z^{k+1}\right).$
@@
Then the lemma is proved by a straightforward computation. ∎
@@ 3.3. Properties of quantum Hamiltonians
Let us prove various properties of quantum Hamiltonian densities $h_{n}$ of dispersionless KdV hierarchy. Our main technique is based on the vertex operator representation of W-type operators and their relations to Hall-Littlewood polynomials, which is studied in [12]. These operators have a deep relationship with and find wide application in the theory of (classical) integrable hierarchy such as KP theory and its reduction (see. e.g., [18] and references therein), and we are now to apply these techniques to study the quantum integrable hierarchy.
@@
The following relation holds true:
@@
Operators $B_{n}$ can be used to define the Schur polynomials ([19]). Given a partition $\lambda=(\lambda_{1},\dots,\lambda_{\ell(\lambda)})$, it follows that
@@
Note that for $\lambda_{1}\geq\lambda_{2}\dots\geq\lambda_{\ell}>0$, the corresponding Hall-Littlewood polynomial coincides with the Schur polynomial.
@@
$=\sum_{a=0}^{n}(-1)^{a-n}\binom{a}{\ell}S_{(a+1,\underbrace{1,\dots,1}_{n-a})},$
@@
then we prove Proposition 1.2, Theorem 1.3, Theorem  1.5 and Proposition 1.6 with $h_{n}$ replaced by $\tilde{h}_{n}$. Finally, we prove that $\tilde{h}_{n}$ actually satisfies the recursion relation (3.1), hence $\tilde{h}_{n}=h_{n}$, and this proves Theorem 3.4.
@@
[…]
@@
(3.9) $\tilde{h}_{n}$ $=\sum_{k=0}^{n}\frac{(-1)^{k}}{k+1}\sum_{a=n-k}^{n}(-1)^{a-n}\binom{a}{n-k}S_{(a+1,\underbrace{1,\dots,1}_{n-a})}$
$=\frac{1}{n+1}\sum_{k=0}^{n}\frac{1}{\binom{n}{k}}S_{(n-k+1,\underbrace{1,\dots,1}_{k})},$
@@
[…]
@@
The actions of $P^{(k)}_{m}$ on Hall-Littlewood polynomials are computed in [18]. Precisely, fix $(\lambda_{1},\dots,\lambda_{\ell})\in\mathbb{Z}^{\ell}$, then
(3.10) $P^{(k)}_{m}(S_{\lambda})=$ $\sum_{i=1}^{\ell}\binom{\lambda_{i}-m-i}{k}S_{(\lambda_{1},\dots,\lambda_{i}-m,\dots,\lambda_{\ell})}+\delta_{m,0}\binom{-\ell}{k+1}S_{\lambda}$
$+\delta_{m<0}\sum_{i=1}^{-m}(-1)^{m-i}\binom{i-\ell-1}{k}S_{(\lambda_{1},\dots,\lambda_{\ell},i,\underbrace{1,\dots,1}_{-m-i})},$
where we set $\delta_{m<0}=1$ for $m<0$ and $\delta_{m<0}=0$ for $m\geq 0$. Let us use this to compute $\mathcal{D}(\tilde{h}_{n})$. Due to (3.9), let us first consider the action
@@
Then it follows that
@@
3.3.3. Proposition 1.5 for $\tilde{h}_{n}$
Let us consider the eigenvalue problems for $\tilde{h}_{n}$. By using (3.10), we see that all the Schur polynomials $S_{\lambda}$ serve as common eigenvectors of $P^{(k)}_{0}$. It follows from (3.8) that
[…]
Then we prove Proposition 1.5 for $\tilde{h}_{n}$ by applying (3.10). Furthermore, we recover the $\hbar$-dependence by noticing that $\varphi(\tilde{h}_{n})_{(n)}$ is of differential degree $-n-1$ and $S_{\lambda}$ is of differential degree zero.
@@
=\,= $\sum_{a,b\in\mathbb{Z}}\binom{a}{m}\binom{b}{n}B^{*}_{a-n}B_{a}B^{*}_{b-m}B_{b}-B^{*}_{b-m}B_{b}B^{*}_{a-n}B_{a}$
=\,= $\sum_{a\in\mathbb{Z}}\binom{a}{n}\binom{a+m}{m}B^{*}_{a-n}B_{a+m}-\sum_{a\in\mathbb{Z}}\binom{a}{n}\binom{a-n}{m}B^{*}_{a-n-m}B_{a}$
=\,= $0,$
@@
(3.11) $[A_{(n)},B_{(m)}]=\sum_{i\geq 0}\binom{n}{i}(A_{(i)}B)_{m+n-i},\quad A,B\in V,\quad n,m\in\mathbb{Z},$

### v3

@@ 2.2. Quantum Hamiltonian structure
where $f,g\in\mathcal{A}[\hbar]$ are arbitrary choices of densities of $F$ and $G$, respectively. This Lie bracket defines a deformation quantization of the Hamiltonian structure (2.6) in the sense that
@@
In particular, we have proved Theorem 1.1 and Proposition 1.4 as special cases of Theorem 2.9 for $N=1$ and $\eta=1$.
As an application of the above formula, let us prove the following interesting fact which will be used later.
@@
The assumption $a_{(0)}=0$ implies that for any $g\in\mathcal{A}$, $\varphi(f)_{(0)}\varphi(g)=0$. Then by taking the $\hbar$-coefficient of $\varphi(f)_{(0)}\varphi(g)$, we arrive at
@@
here we denote by $\eta_{\alpha\beta}$ the inverse of $\eta^{\alpha\beta}$. Due to (2.13), we see that $A=0$, therefore we conclude that
@@
which implies that actually $\varphi(f_{0})_{(0)}=0$. Hence, we replace $f$ by $(f-f_{0})/\hbar$, and the same argument shows that $f_{1}$ is of the same form as $f_{0}$. In this way, we prove that actually each $f_{n}$ has the same form of $f_{0}$, namely, there exist polynomials $B_{\alpha}(\hbar),C(\hbar)\in\mathbb{C}[\hbar]$ such that
@@ 3. Weyl quantization of dispersionless KdV hierarchy
$H_{2}^{c},H_{3}^{c}]=\int\frac{\hbar}{4}v^{4}v_{x}+\frac{\hbar^{2}}{12}(3vv_{x}v_{xx}+v^{2}v_{xxx})+\frac{\hbar^{3}}{720}v^{(5)}\neq 0.$
The quantization problem is then to find quantum Hamiltonians $H_{n}\in\mathcal{F}[\hbar]$ such that
$H_{n},H_{m}]=0,\quad\lim_{\hbar\to 0}H_{n}=H_{n}^{c}.$
@@ 3.1. Construction of quantum Hamiltonians
From now on, we will consider the Heisenberg vertex algebra of rank 1 with $\eta=1$. Our idea to construct the quantum Hamiltonians is as follows. The ring $\mathcal{A}$ of differential polynomials is a commutative algebra, and we can view the normal order product on $V$ as a deformation of this algebra structure, in the sense that
@@
which follows from Corollary  2.8. The normal order product is neither commutative nor associative, and we can use the idea of Weyl quantization in quantum mechanics to quantize the Hamiltonians $H_{n}^{c}$. Roughly speaking, in Weyl quantization, a quantum Hamiltonian is obtained from a classical Hamiltonian by averaging all the possible ways to compose all the quantum operators appearing in Hamiltonian. For example, let $(x,p)$ be a pair of conjugated variables, then we have the following Weyl quantization:
[…]
@@
To give a precise definition of the non-associative Weyl quantization of Hamiltonians, let us define a set $BT(n)$ of (n1)!(n-1)! elements for all $n\geq 1$. Each element in $BT(n)$ is a full binary tree whose nodes, except for the root node, are labeled either by $L$ (for left) or $R$ (for right) and the root node is labeled by an integer $i$ with $1\leq i\leq(n-1)!$. In addition, all $n$ leaves of a tree in $BT(n)$ are labeled by $1,\dots,n$.
We proceed to define $BT(n)$ recursively. First, define $BT(1)$ to be the set $\{1\}$, where we interpret the element $1\in BT(1)$ as a single root node labeled by 1. Define $BT(2)$ to be the set of the unique full binary tree with two leaves labeled by $L,1$ and $R,2$ and the root labeled by 1.
1L,1R,2
@@
To define the set $BT(n+1)$ for $n\geq 2$, we attach a leaf node to trees in $BT(n)$ and change the labels accordingly. Take a tree $T\in BT(n)$ whose root node is labeled by $k$ and fix a choice of its leaf with label $i\in\{1,\dots,n\}$, let us define a new tree $T_{i}$ as follows. Firstly, we make the chosen leaf with label $i$ an internal node by attaching to it 2 leaves, and hence obtain a full binary tree with $n+1$ leaves. Then the label ‘$i$’ of this node is removed. For leaves originally labeled by $1,\dots,i-1$ their labels remain unchanged, and for leaves originally labeled by $i+1,\dots,n$ their labels are increased by one. Finally, the newly added leaves are labeled as $L,i$ and $R,i+1$, and the label of the root is changed from $k$ to (k1)n+i(k-1)n+i. We then define
@@
As examples, the elements in $BT(3)$ are presented in Fig. 2 and the elements in $BT(4)$ are presented in Fig. 4 and Fig. 5.
1LL,1R,2R,3
2L,1RL,2R,3
@@
Remark 3.1.
In $BT(n)$, there are distinct elements with the same binary tree structure. For example, we see that in $BT(4)$, there are only 5 distinct trees. However, there are 6 distinct elements in $BT(4)$ as labeled trees. This is also the reason why $BT(n)$ consists of (n1)!(n-1)! elements rather than just $C_{n}$ elements, where $C_{n}$ is the $n$-th Catalan number counting the number of full binary trees with $n$ leaves.
Let us proceed to associate a state $A_{T}$ to each labeled tree $T\in BT(n)$. For $1\in BT(1)$, we simply set $A_{1}=b$. For $n\geq 2$ and $T\in BT(n)$, we first place the state $b$ on every leaf, and then recursively place the state at all other nodes obtained by taking the normal order product of states associated with its left child and right child. More precisely, given a node which is not a leaf, assume that we have put $f\in V$ at its left child and $g\in V$ at its right child, then this node is associated with the state $fg:\in V$. We denote by $A_{T}\in V$ the state associated with the root of $T$. For example, the state associated to the unique tree in $BT(2)$ is :$b$::$b$:, and for $T_{1},T_{2}\in BT(3)$ as presented in Fig. 2, we have
@@
Proof.
We observe that each tree in $BT(n)$ can be decomposed into two subtrees by removing the root node. Then this lemma is verified by a straightforward computation. ∎
By using the above notation, we define differential polynomials
$h_{n}=\frac{1}{(n+1)!}\varphi^{-1}(Z_{n+1}),\quad n\geq 0.$
It follows from Corollary 2.8 that
@@
therefore the Hamiltonians defined by $H_{n}=\int h_{n}$ is a deformation of classical Hamiltonians $H_{n}^{c}$ in the sense that
$\lim_{\hbar\to 0}H_{n}=H_{n}^{c}.$
By replacing $Z_{i}$ by $i!\varphi(h_{i-1})$ in Lemma  3.1, it follows that we can equivalently define $h_{n}$ recursively by
@@
It is easy to verify that $h_{0}=v$ and $h_{1}=\frac{v^{2}}{2}$. Let us compute $h_{2}$. First it follows from the definition of normal order product that
$b:bb::=b_{(-1)}b_{(-1)}b_{(-1)}|0\rangle,$
and by using the quasi-associativity (2.2), we have
$bb:b:=:b:bb::+:\left(\int_{0}^{T}d\lambda\,b\right)[b_{\lambda}b]:=:b:bb::+\hbar T^{2}b,$
@@
$b_{\lambda}b]=\sum_{n\geq 0}\frac{\lambda^{n}}{n!}b_{(n)}b_{(-1)}|0\rangle=\hbar\lambda|0\rangle.$
@@
Let us compute $h_{3}$. We begin by enumerate elements in $BT(4)$. Denote by $R$ and $S$ the first and the second tree presented in Fig. 2, then by definition we arrive at
@@
These trees are presented in Fig. 4 and Fig. 5.
1LLL,1R,2R,3R,4
2LL,1RL,2R,3R,4
3LL,1R,2RL,3R,4
@@
Let us compute the contribution from each graph. First we note that
@@
$b^{2}b^{2}:-b^{4}=2\left(\int_{0}^{T}d\lambda\,b\right)[b_{\lambda}b^{2}]=2\hbar:T^{2}bb:,\quad b^{4}=\,:b:b:bb:::,$
@@
$b_{\lambda}b^{2}]=:[b_{\lambda}b]b:+:b[b_{\lambda}b]:+\int_{0}^{\lambda}[[b_{\lambda}b]_{\mu}b]d\mu=2\hbar\lambda b.$
From the above computation we arrive at
@@
$\left(-2\hbar\frac{\partial}{\partial\hbar}+\sum_{k\geq 0}kv^{(k)}\frac{\partial}{\partial v^{(k)}}\right)h_{n}$ $=0,\quad n\geq 0.$
@@
It follows from the definition and Lemma 2.4 that $Z_{n}$ is of conformal weight $n$. The conformal degree of $V$ induces a gradation on $\mathcal{A}[\hbar]$ via $\varphi$, with $v^{(k)}$ of degree $k+1$, hence we prove the first identity.
@@
This gradation induces a gradation on $V$, where $\deg_{\partial_{x}}b_{(n)}=-n-1$. For $f\in\mathcal{A}[\hbar]$ of differential degree $d$, it is easy to see, using Proposition 2.2 that
@@
Then we can prove by induction, using recursion (3.1), that each $h_{n}$ is of differential degree 0. The lemma is proved. ∎
@@ 3.2. Another expression for quantum Hamiltonians
As we have already seen, it is not easy to compute each $h_{n}$ using the definition. In what follows, we give another expression of $h_{n}$, which simplifies computations and more importantly, enables us to prove various properties of quantum Hamiltonians. From now on, for simplicity, we will set $\hbar=1$, and the dependence of $\hbar$ can be recovered from homogeneous conditions.
@@
This recursion relation makes it easy to compute each $g_{n}$, and in particular, it follows that $g_{n}$ are given by Schur polynomials. Recall that for an integer partition $\lambda=(\lambda_{1},\dots,\lambda_{\ell(\lambda)})$, we associate it with a differential polynomial $S_{\lambda}\in\mathcal{A}$, called the Schur polynomial associated with the partition $\lambda$, by
@@
the generating function of $g_{n}$, then it follows from the recursion relation of $g_{n}$ that this function is the unique formal series solution of the equation
@@
Therefore, it suffices to show that the formal power series
$\frac{W(z)}{z}-\frac{1}{z}$
solves the above equation, with
$W(z)=\exp\left(\sum_{k\geq 0}\frac{v^{(k)}}{(k+1)!}z^{k+1}\right).$
@@
Then the lemma is proved by a straightforward verification. ∎
@@ 3.3. Properties of quantum Hamiltonians
Let us prove various properties of quantum Hamiltonian densities $h_{n}$ of quantum dispersionless KdV hierarchy. Our main technique is based on the vertex operator representation of W-type operators and their relations to Hall-Littlewood polynomials, which is studied in [13]. These operators have a deep relationship with and find wide application in the theory of (classical) integrable hierarchies such as KP theory and its reduction (see, e.g., [19] and references therein), and we are now to apply these techniques to study the quantum integrable hierarchy.
@@
The following relation holds true for any $k\geq 0$ and $m\in\mathbb{Z}$:
@@
Operators $B_{n}$ can be used to define the Schur polynomials ([20]). Given an integer partition $\lambda=(\lambda_{1},\dots,\lambda_{\ell(\lambda)})$, it follows that
@@
It follows that for $\lambda_{1}\geq\lambda_{2}\dots\geq\lambda_{\ell}>0$, the corresponding Hall-Littlewood polynomial coincides with the Schur polynomial.
@@
(3.6) $=\sum_{a=0}^{n}(-1)^{a-n}\binom{a}{\ell}S_{(a+1,\underbrace{1,\dots,1}_{n-a})},$
@@
then we prove Proposition 1.2, Theorem 1.3, Theorem 1.5 and Proposition 1.6 with $h_{n}$ replaced by $\tilde{h}_{n}$. Finally, we prove that $\tilde{h}_{n}$ actually satisfies the recursion relation (3.1), hence $\tilde{h}_{n}=h_{n}$, and this proves Theorem 3.4.
@@
(3.8) $\varphi^{-1}\mathchoice{\mathrel{\raisebox{0.86108pt}{$\scriptstyle\circ$}}}{\mathrel{\raisebox{0.86108pt}{$\scriptstyle\circ$}}}{\mathrel{\raisebox{0.43057pt}{$\scriptscriptstyle\circ$}}}{\mathrel{\raisebox{0.43057pt}{$\scriptscriptstyle\circ$}}}Y(\varphi(\tilde{h}_{n}),z)\mathchoice{\mathrel{\raisebox{0.86108pt}{$\scriptstyle\circ$}}}{\mathrel{\raisebox{0.86108pt}{$\scriptstyle\circ$}}}{\mathrel{\raisebox{0.43057pt}{$\scriptscriptstyle\circ$}}}{\mathrel{\raisebox{0.43057pt}{$\scriptscriptstyle\circ$}}}\varphi=\sum_{k=0}^{n}\frac{(-1)^{n-k}}{(n-k+1)!}\partial_{z}^{n-k}P^{(k)}(z),$
@@
$\tilde{h}_{n}$ $=\sum_{k=0}^{n}\frac{(-1)^{k}}{k+1}\sum_{a=n-k}^{n}(-1)^{a-n}\binom{a}{n-k}S_{(a+1,\underbrace{1,\dots,1}_{n-a})}$
(3.9) $=\frac{1}{n+1}\sum_{k=0}^{n}\frac{1}{\binom{n}{k}}S_{(n-k+1,\underbrace{1,\dots,1}_{k})},$
@@
$\mathcal{D}=2\,\varphi^{-1}\mathchoice{\mathrel{\raisebox{0.86108pt}{$\scriptstyle\circ$}}}{\mathrel{\raisebox{0.86108pt}{$\scriptstyle\circ$}}}{\mathrel{\raisebox{0.43057pt}{$\scriptscriptstyle\circ$}}}{\mathrel{\raisebox{0.43057pt}{$\scriptscriptstyle\circ$}}}\varphi(\tilde{h}_{2})_{(1)}\mathchoice{\mathrel{\raisebox{0.86108pt}{$\scriptstyle\circ$}}}{\mathrel{\raisebox{0.86108pt}{$\scriptstyle\circ$}}}{\mathrel{\raisebox{0.43057pt}{$\scriptscriptstyle\circ$}}}{\mathrel{\raisebox{0.43057pt}{$\scriptscriptstyle\circ$}}}\varphi\in\mathrm{End}(\mathcal{A}),$
@@
The actions of $P^{(k)}_{m}$ on Hall-Littlewood polynomials are computed in [19]. Specifically, fix $(\lambda_{1},\dots,\lambda_{\ell})\in\mathbb{Z}^{\ell}$, then for any $k\geq 0$ and $m\in\mathbb{Z}$, we have
$P^{(k)}_{m}(S_{\lambda})=$ $\sum_{i=1}^{\ell}\binom{\lambda_{i}-m-i}{k}S_{(\lambda_{1},\dots,\lambda_{i}-m,\dots,\lambda_{\ell})}+\delta_{m,0}\binom{-\ell}{k+1}S_{\lambda}$
(3.10) $+\delta_{m<0}\sum_{i=1}^{-m}(-1)^{m-i}\binom{i-\ell-1}{k}S_{(\lambda_{1},\dots,\lambda_{\ell},i,\underbrace{1,\dots,1}_{-m-i})},$
where we set $\delta_{m<0}=1$ for $m<0$ and $\delta_{m<0}=0$ for $m\geq 0$. Let us use this to compute $\mathcal{D}(\tilde{h}_{n})$. Due to (3.9), we first consider the action
@@
We arrive at the following computation:
@@
3.3.3. Theorem 1.5 for $\tilde{h}_{n}$
Before we prove Theorem 1.5 for $\tilde{h}_{n}$, let us explain why it is necessary to take $\varphi(h_{n})_{(n)}\in\mathrm{End}(V)$ for studying the eigenvalue problem. We have proved in Lemma 3.2 that $\varphi(h_{n})\in V$ is of conformal weight $n+1$, and it follows from Lemma 2.4 that the operator $\varphi(h_{n})_{(m)}$ is of conformal degree nmn-m. It follows that $\varphi(h_{n})_{(m)}$ admits non-zero eigenvectors unless $m=n$.
Let us proceed to study the eigenvalue problems for $\tilde{h}_{n}$. By using (3.10), we see that all the Schur polynomials $S_{\lambda}$ serve as common eigenvectors of $P^{(k)}_{0}$. It follows from (3.8) that
$\varphi^{-1}\mathchoice{\mathrel{\raisebox{0.86108pt}{$\scriptstyle\circ$}}}{\mathrel{\raisebox{0.86108pt}{$\scriptstyle\circ$}}}{\mathrel{\raisebox{0.43057pt}{$\scriptscriptstyle\circ$}}}{\mathrel{\raisebox{0.43057pt}{$\scriptscriptstyle\circ$}}}\varphi(\tilde{h}_{n})_{(n)}\mathchoice{\mathrel{\raisebox{0.86108pt}{$\scriptstyle\circ$}}}{\mathrel{\raisebox{0.86108pt}{$\scriptstyle\circ$}}}{\mathrel{\raisebox{0.43057pt}{$\scriptscriptstyle\circ$}}}{\mathrel{\raisebox{0.43057pt}{$\scriptscriptstyle\circ$}}}\varphi=\frac{1}{n+1}\sum_{k=0}^{n}\binom{n+1}{k}P^{(k)}_{0}$.
Then we prove Proposition 1.5 for $\tilde{h}_{n}$ by applying (3.10). Furthermore, we recover the $\hbar$-dependence by noticing that $\varphi(\tilde{h}_{n})_{(n)}$ is of differential degree n1-n-1 and $S_{\lambda}$ is of differential degree zero.
@@
$=\,$ $\sum_{a,b\in\mathbb{Z}}\binom{a}{m}\binom{b}{n}B^{*}_{a-n}B_{a}B^{*}_{b-m}B_{b}-B^{*}_{b-m}B_{b}B^{*}_{a-n}B_{a}$
$=\,$ $\sum_{a\in\mathbb{Z}}\binom{a}{n}\binom{a+m}{m}B^{*}_{a-n}B_{a+m}-\sum_{a\in\mathbb{Z}}\binom{a}{n}\binom{a-n}{m}B^{*}_{a-n-m}B_{a}$
$=\,$ $0,$
@@
(3.11) $\left[A_{(n)},B_{(m)}\right]=\sum_{i\geq 0}\binom{n}{i}(A_{(i)}B)_{(m+n-i)},\quad A,B\in V,\quad n,m\in\mathbb{Z},$